OpenStudy (mackenzie2013):

Perform the indicated operation. 6 1/3 + 8 3/4 A-13 1/4 B- 14 1/4 C- 15 1/12 D- 15 3/7

OpenStudy (wolf1728):

6 1/3 + 8 3/4 You need to find the lowest common denominator which is 12 So the fractions become 6 4/12 + 8 9/12 = 14 13/12 or 15 1/12 answer is C 15 1/12

OpenStudy (mackenzie2013):

Thank you! :)

OpenStudy (anonymous):

easy anymore

OpenStudy (anonymous):

19/3 + 35/4 = ((19 × 4) + (35 × 3)) / (3 × 4) = (76 + 105) / 12 = 181/12 = 181/12 = 15 1/12

OpenStudy (anonymous):

anymore of those plz say there is more!!!!!!!!

OpenStudy (mackenzie2013):

It was C :) thanks

OpenStudy (anonymous):

anymore